OpenStudy (anonymous):

-> Need help <- Question 1 (Multiple Choice Worth 2 points) A cab charges $1.45 for the first mile and $0.55 for each additional mile. Write and solve an inequality to determine how many miles Ariel can travel if she has $35 to spend. $1.45 + $0.55x ≥ $35; x ≥ 61 miles $1.45 + $0.55x ≤ $35; x ≤ 61 miles $0.55 + $1.45x ≥ $35; x ≥ 23 miles $0.55 + $1.45x ≤ $35; x ≤ 23 miles ________________________________________

OpenStudy (anonymous):

1.45 + (0.55)(x-1) <= 35 0.90 +0.55 x <= 35 But it is not one of your options???

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

how did you figure this out? could you explain this to me? :)

OpenStudy (anonymous):

Where am I going wrong? first mile is !.45, 2nd one, the next one is 1.45 + (2-1)0.55=2.00 = 0.90 + 2 x 0.55 = 2.00

OpenStudy (anonymous):

I never said you were wrong :) I just wanted to know how you solved this so I can understand :)

OpenStudy (anonymous):

No, I was asking that of myself. First mile is 1.45. The rest are 1.45 plus 0.55 for every mile beyond the fits one, thus, 0.55(x-1) miles.

OpenStudy (anonymous):

oh okay! lol well thank you for the help :)

OpenStudy (anonymous):

And where do the inequalities for x by itself come from? Presumably from solving the inequalities for the cost. So rule out "x>=" options as limit must be "x<=".